Hi The data from 2A1723 was found by Kris to look suspicious in the RICH. Since then we identified the problem as being bad calibration of RICH pedestal, in so fra the pedestal changed between the pedestal run and the first datarun for 15240 causing spurious hits in the RICH. This change was apperently permanent thorughout the run5 pp. New gtrs and dst has been produced (15240-15249). Flemming -------------------------------------------- Flemming Videbaek Physics Department Bldg 510-D Brookhaven National Laboratory Upton, NY11973 phone: 631-344-4106 cell: 631-681-1596 fax: 631-344-1334 e-mail: videbaek @ bnl gov _______________________________________________ Brahms-dev-l mailing list Brahms-dev-l_at_lists.bnl.gov https://lists.bnl.gov/mailman/listinfo/brahms-dev-lReceived on Fri Jun 29 2007 - 14:22:51 EDT
This archive was generated by hypermail 2.2.0 : Fri Jun 29 2007 - 14:23:18 EDT